OpenStudy (anonymous):

A yearbook printer charges based on the number of pages printed. Here is a table that shows the cost of some recent yearbooks. Number of pages 50 100 150 200 Cost 1.55 2.25 2.95 3.65 Which equation gives the cost, y, in terms of the number of pages, x? A. y = 0.014x + 49.15 B. y = 0.014x + 1.55 C. y = 0.014x + 0.85 D. y = 58.8x + 0.85 can someone please help and fast

OpenStudy (cal.lavender):

First, you have to find the slope. (3.65 -2.95 )/ (200-150) =m And then you use like the point-slop form, you can just take a random point and plug it into the equation and solve. For example, y-3.65= 0.014 (x-200) and you distribute 0.013x-2.8=y-3.65 Add 3.65 on both sides y=0.014x +0.85 Sorry if this was confusing.. D:

OpenStudy (anonymous):

thanks